Circular No. 5891

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET MUELLER (1993s) Jean Mueller reports her discovery of a comet on a plate taken with

the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. The object is diffuse, of mag 17.5-18, with a pronounced tail ∼ 80′′ long toward p.a. 270^{o}. Confirmation was obtained via a second plate taken by J. D. Mendenhall and Mueller. The following positions are measurements by Mueller of the centers of the short, diffuse trails:

1993 UT  α_{2000}  δ_{2000}
Nov. 20.41910  6^{h}58^{m}55.^{s}65  +20^{o}35′20′′.0
21.36111  6 58 42.36  +20 37 31.9

RX J2107.9–0518 N. M. Shakhovskoj, Crimean Astrophysical Observatory; and S. V.

Kolesnikov, Odessa University, report the detection of circular polarization of the eclipsing polar RX J2107.9–0518. The observation was made with the 2.6-m Shajn telescope, using a single-channel polarimeter with a fast-rotating half-wave plate in the broad R band (500-750 nm) and for a 4-s integration time. The circular polarization reached its maximum positive value of 3.08 ± 0.85 percent on Aug. 15.955 UT (phase 0.29) and its maximum negative value of −8.80 ± 1.10 percent on Aug. 16.002 (phase 0.83). These data are the means from a series of 64 integrations. Simultaneous UBVRI photometry with the 1.25-m telescope gave the following mean brightness and colors outside eclipse: V = 15.0, U–B = −0.85, B–V = +0.25, V–R = +0.35, R–I = +0.30. The times of the eclipses were very close to the ephemeris by Schwope et al. (1993, A.Ap. 271, L25-L28).

NOVA SAGITTARII 1993 A. C. Gilmore, Mount John University Observatory, reports the follow-

ing photometry made with Johnson-Cousins UBVR_{c}I_{c} filters; the measures were made differentially from SAO 186445 (assumed V = 7.747, U–B = −0.080, B–V = +0.079, V–R_{c} = +0.055, V–I_{c} = +0.16 on Sept. 17.4 UT, and located 10′.7 west and 5′.2 south of N Sgr 1993): Oct. 24.43, V = 11.57 ± 0.02, U–B = −0.43 ± 0.01, B–V = +0.06 ± 0.04, V–R_{c} = +1.31 ± 0.01, V–I_{c} = +1.4 ± 0.3; 25.42, 11.67 ± 0.01, −0.45 ± 0.04, +0.16 ± 0.01, +1.35 ± 0.02, +1.1 ± 0.1; 28.40, 11.94 ± 0.03, −0.55 ± 0.05, +0.16 ± 0.02, +1.56 ± 0.01, –. Conditions were marginal on Oct. 25.

1993 November 21 Daniel W. E. Green